# Re2LLM: Reflective Reinforcement Large Language Model for Session-based Recommendation

### v2

@@ 1. Introduction
Figure 1. Our method (Fig. c) obtains effective task-specific feedback compared with prompt-based (Fig. a) and fine-tuning-based (Fig. b) methods.
@@
  • Reflective Exploration Module leverages the self-reflection ability of LLMs (Madaan et al., 2023; Pan et al., 2023) to extract specialized knowledge that is comprehensible and digestible by LLMs for SBR. Specifically, we employ LLMs to identify common errors in their responses, and then generate corresponding specialized knowledge (i.e., hints) to rectify these errors through LLMs’ self-reflections. Hence, the specialized knowledge can align seamlessly with LLM comprehension as it is summarized by the LLM itself. Besides, we construct a hint knowledge base that serves as a pool to maintain specialized knowledge (hints), using an automated filtering strategy to ensure the effectiveness and diversity of retained hints.
@@ 2.3. Self-reflection in Large Language Models
Recent advances in prompting strategies have effectively enhanced LLMs’ ability to handle complex tasks. The Chain-of-Thought principle (Wei et al., 2022) allows LLMs to reason through a specified path. Similarly, the Tree-of-Thought strategy (Yao et al., 2023b) aids in exploring multiple reasoning paths and self-assessing various scenarios. Despite their impressive performance, LLMs are still prone to hallucination and incorrect reasoning (Pan et al., 2023). Consequently, much effort is devoted to exploring LLMs’ self-reflection (Pryzant et al., 2023; Madaan et al., 2023; Yao et al., 2023a; Shinn et al., 2023) capabilities, where they iteratively refine outputs based on self-generated reviews or feedback without additional training. For example, Madaan et al. (2023) apply this feedback-and-refine framework to improve LLMs’ performance across NLP tasks. Pryzant et al. (2023) improve prompts by having an LLM critique on generations and then refine them based on its feedback. Yao et al. (2023a) train a retrospective language model as a prompt generator, using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) based on environment-specific rewards for optimization. However, the potential of LLMs’ self-reflection in recommendation remains underexplored. To the best of our knowledge, only DRDT (Wang et al., 2023) is related to reflection for recommendation, which triggers LLMs to conduct the iterative reflection on a specific session until hitting the ground truth item. However, it is essentially a case-by-case reflection strategy, which hinders summarising the specialized knowledge from global sessions. To this end, we propose to explore different sessions to maintain a hint knowledge base, whose knowledge (i.e., hints) can be utilized to enhance LLMs reasoning for all sessions.
@@ 3. Methodology
Figure 2. The overall architecture of our proposed Re2LLM approach.
@@ 3.1.1. Multi-Round Self-Reflection Generation
To activate the self-reflection ability of LLMs, we propose to identify errors in LLMs’ inference by comparing their predictions against the ground truth. Specifically, we first instruct the LLM to generate a top-10 ranking list $O(s)$ from a candidate set $\mathcal{C}$ containing $|\mathcal{C}|$ items based on current session $s$:
@@ 2. Popularity Bias: …
Then, we identify the incorrectly predicted sessions where the top-10 ranking list fails to hit the ground truth item at time step $l+1$ (i.e., $v_{l+1}^{s}\notin O(s)\O$). Afterward, we prompt the LLM to conduct self-reflection, that is, analyze and summarize the prediction errors, to generate potential hints in natural language to rectify such errors. These hints thus can be considered as specialized knowledge for better recommendations. Inspired by Chain-of-Thought prompting (Wei et al., 2022), our approach focuses on analyzing and alleviating these errors in a step-by-step and progressive reasoning chain, producing hints that are tailored to be compatible with LLM comprehension capabilities, as introduced below.
@@ 3.1.2. Automated Filtering
To ensure the effectiveness of hints in $\mathcal{H}$, we incorporate a new hint only if it can benefit performance improvement for recommendation. Specifically, we first construct a hint-enhanced prompt $P^{*}$ (Prompt 5) to trigger the LLM with the hint-enhanced prompt for recommendation inference:
(2) $O^{*}(s)=LLM(P^{*}(s,\mathcal{C},h)),$
where $O^{*}(s)$ is the recommendation list for session $s$ based on the hint-enhanced prompt $P^{*}$. Then, we compare $O^{*}(s)$ with $O(s)$ obtained by Eq. 1 which is obtained based on the basic prompt without hint enhancement. The hint $h$ can be recognized as effective if it leads to a better prediction by the LLM, i.e., $v_{l+1}^{s}\in O^{*}(s)\And v_{l+1}^{s}\notin O(s)$.
@@
Meanwhile, some hints may be redundant generated by LLMs, as multiple sessions may suffer from similar causes. For example, ‘Consider the release years of movies in the user’s history for era preference’ and ‘Think about the production year of watched movies, recommend movies in that ear from the candidate set’ share a similar semantic meaning. The redundancy of hints may result in a high cost of the maintenance and utilization of the hint knowledge base $\mathcal{H}$, e.g., larger size of the hint knowledge base and increased complexity in retrieving the targeted hint. To reduce the redundancy in the hint knowledge base $\mathcal{H}$, we only incorporate new hints that are distinct from existing ones. Specifically, we employ LLMs to detect the semantic similarity between the candidate hint $h^{\prime}$ and existing ones $\{h\in\mathcal{H}\}$, given by,
@@ 3.2. Reinforcement Utilization Module
To guide LLMs to infer more accurate SBR predictions, we propose to utilize the constructed hint knowledge base to prevent errors in the inference process of LLMs. In practical scenarios, the absence of explicit labels on the hints’ efficacy leads to a challenge in conducting supervision on hint selection. To this end, we employ the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm in DRL to ensure stable and efficient training of our retrieval agent. Specifically, our agent is trained to select the most relevant hints based on session-related context information, thereby preventing LLMs from similar reasoning mistakes for recommendation.
@@ 3.2.1. Markov Decision Process (MDP)
Transition. In each step, the agent can receive task-specific feedback by simulating the real-world RSs. The agent observes the state of the current session, takes action to select a hint for the LLM, and receives a reward from the environment. Subsequently, the environment transits to the next session’s state, the transition is denoted as $T(\mathbf{z})=\mathbf{z}^{\prime}$. This setup can flexibly accommodate the scenario where multiple steps per session are involved for more sophisticated learning processes in our future studies.
Replay buffer. To facilitate efficiency in policy optimization, we maintain a replay buffer $D=(\mathbf{z},\mathbf{a},r,\mathbf{z}^{\prime})$ to store the tuples of observation state, agent action, reward, and next observation state. With the records in the replay buffer, the retrieval agent can refine successful strategies and learn from erroneous trials.
@@ 3.2.2. PPO Training
where the first term is the reward measured by recommendation tasks, and the second term is the KL-divergence of policies with a coefficient $\beta$ for the regulation in policy updates.
In summary, we train a retrieval agent with task-specific feedback via DRL, which can select the appropriate action to improve LLM’s performance through hint-enhanced prompts for better recommendations.
3.2.3. Retrieval-Enhanced LLMs for Recommendation
With the constructed knowledge base and the trained retrieval agent, we achieve more accurate recommendations by carefully selecting appropriate hints for prompt enhancement during inference. The final recommendation output $\hat{O}$ can be denoted as follows:
(6) $\hat{O}=LLM(P^{*}(s,\mathcal{C},\pi_{\bm{\theta}}(\mathbf{z}_{s}))),$
Input: Training data $\mathcal{S}_{train}$ for SBR, hint knowledge base size $n_{h}$.
@@ 3.2.3. Retrieval-Enhanced LLMs for Recommendation
Input: Training data $\mathcal{S}_{train}$ for SBR, constructed $\mathcal{H}$, max episode $N$.
@@ 4.1.2. Baselines
We compare our proposed method with eight state-of-the-art baseline methods [^5] Our method does not require fine-tuning, and due to the high computational cost, fine-tuning-based methods are not in our scope. , and the details are listed as follows. FPMC (Rendle et al., 2010) combines matrix factorization with the first-order Markov chain. GRU4Rec (Hidasi et al., [n. d.]) uses Gated Recurrent Unit (GRU), a typical RNN layer, to model whole sessions. NARM (Li et al., 2017) introduces the attention mechanism into RNNs to better model the sequential behavior. SRGNN (Wu et al., 2019) constructs session graphs and employs GNN layers to obtain embeddings. GCEGNN (Wang et al., 2020) further includes global-level information as an additional graph to enhance the model performance. AttenMixer (Zhang et al., 2023b) achieves multi-level reasoning over item transitions by the attention-based readout method. LLMRank (Hou et al., 2023) demonstrates the promising zero-shot inference ability of LLMs through recency-focused prompting and in-context learning. NIR (Wang and Lim, 2023) employs a multi-step prompt to capture user preferences first, then select representative movies, and finally perform the next-item prediction.
@@
4.1.3. Evaluation Strategy and Metrics
For each dataset, we apply the split-by-ratio strategy (Sun et al., 2020) for sessions to obtain training, validation, and test sets with a ratio of 7:1:2. For the full dataset setting, we use the entire training set. For the few-shot setting, we sample 500 training samples from the entire training set. In terms of evaluation, we adopt a leave-one-out evaluation methodology, which considers the last item in each session as the target. Considering the cost and efficiency of LLMs, we randomly sample a subset of 2,000 sessions from the test set. For the efficiency of evaluation, we sampled 49 negative items for each positive item (i.e., target). In addition, we adopt two commonly used metrics for model evaluation, including normalized discounted cumulative gain ($NDCG@n$) and hit rate ($HR@n$), with $n=\{5,10\}$. The results show the average of five runs with different random seeds.
4.1.4. Implementation Details
For a fair comparison, all methods are optimized by the Adam optimizer with the same embedding dimension $768$ that aligns with the BERT encoder dimension [^7] Empirically, we discover a consistent performance drop for all methods if using MLP to reduce dimensionality here.. Following (Sun et al., 2022), we employ the Optuna [^8] https://optuna.org/ framework to efficiently optimize the hyper-parameter of all methods. We conduct 20 trials on the following searching spaces, i.e., learning rate: {$1e^{-2},1e^{-3},1e^{-4}$}, weight decay: {$1e^{-4},1e^{-6}$ $,1e^{-8}$}, and batch size: {$16,64,256$}. For our method Re2LLM, we set candidate set size $|\mathcal{C}|$ to 50, knowledge base size $n_{h}$ to 20, and few-shot training size $n_{e}$ to 500. The impacts of essential hyper-parameters on Re2LLM can be found in Section 4.5. For LLM-based methods, we use ‘gpt-4’ API as the backbone model for these methods. For the rest hyper-parameters in baseline methods, we follow the suggested settings in the original papers. The experiments are run on a single Tesla V100-PCIE-32GB GPU.
@@ 4.3. Albation Studies (RQ2)
Table 3 shows the performance of all ablation studies. First, Re2LLM outperforms its variant w/o-HE, which indicates the necessity of activating LLM’s reflection mechanism with hint enhancement in our Reflective Exploration Module. Second, Re2LLM outperforms variants w/o-DRL(RAN) and w/o-DRL(ALL), indicating the importance of our Reinforcement Utilization Module for the training of the retrieval agent with task-specific feedback signals. The variant w/o-DRL(RAN) has the worst performance by random sampling because inaccurate hints can mislead the LLM inference for recommendation, resulting in negative impacts. The variant w/o-DRL(ALL) shows limited improvement compared to the variant w/o-HE, revealing that simply feeding all hints into LLMs without session-aware selection is not optimal. Third, the variant w/o-AF achieves relatively better performance among all variants. Nevertheless, it still underperforms our Re2LLM due to the positive impact of the automated filtering mechanism on the hint knowledge base construction. In summary, the above results exhibit the efficacy of different modules in the proposed Re2LLM for more accurate SBR.

### v4

@@ 1. Introduction
Figure 1. Our method (c) obtains effective task-specific feedback compared with prompt-based (a) and fine-tuning-based (b) methods.
@@
  • Reflective Exploration Module leverages the self-reflection of LLMs (Madaan et al., 2023; Pan et al., 2023) to extract specialized knowledge that is comprehensible and digestible by LLMs for SBR. Specifically, we employ LLMs to identify common errors in their responses, and then generate corresponding specialized knowledge (i.e., hints) to rectify these errors through LLMs’ self-reflections. Hence, the specialized knowledge can align seamlessly with LLM comprehension as it is summarized by the LLM itself. Besides, we construct a hint knowledge base that serves as a pool to maintain specialized knowledge (hints), using an automated filtering strategy to ensure the effectiveness and diversity of retained hints.
@@ 2.3. Self-reflection in Large Language Models
Recent advances in prompting strategies have effectively enhanced LLMs’ ability to handle complex tasks. Chain-of-Thought (Wei et al., 2022) and Tree-of-Thought (Yao et al., 2023b) strategies allow LLMs to reason through a specified path. However, LLMs are still prone to hallucination and incorrect reasoning (Pan et al., 2023). Consequently, much effort is devoted to exploring LLMs’ self-reflection (Pryzant et al., 2023; Madaan et al., 2023; Yao et al., 2023a; Shinn et al., 2023) capabilities, where they iteratively refine outputs based on self-generated reviews or feedback without additional training. For example, Madaan et al. (2023) apply this feedback-and-refine framework to improve LLMs’ performance across NLP tasks. Pryzant et al. (2023) improve prompts by having an LLM critique on generations and then refine them based on its feedback. Yao et al. (2023a) train a retrospective language model as a prompt generator, using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) based on environment-specific rewards for optimization. However, the potential of LLMs’ self-reflection in recommendation remains underexplored. To the best of our knowledge, only DRDT (Wang et al., 2023) is related to reflection for recommendation, which triggers LLMs to conduct the iterative reflection on a specific session until hitting the ground truth item. However, it is essentially a case-by-case reflection strategy, which hinders summarising the specialized knowledge from global sessions. To this end, we propose to explore different sessions to maintain a hint knowledge base, whose knowledge (i.e., hints) can be utilized to enhance LLMs reasoning for all sessions.
@@ 3. Methodology
Figure 2. The overall architecture of our proposed Re2LLM approach. The overall flowchart (green) contains a Reflective Exploration Module (yellow) for the generation of session-aware hints as specialized knowledge and a Reinforcement Utilization Module (blue) for the learning to retrieve obtained specialized knowledge.
@@ 3.1.1. Multi-Round Self-Reflection Generation
To activate the self-reflection ability of LLMs, we propose to identify errors in LLMs’ inference by comparing their predictions against the ground truth. Specifically, we first instruct the LLM to generate a ranking list $O(s)$ from a candidate set $\mathcal{C}$ containing $|\mathcal{C}|$ items based on current session $s$:
@@
Then, we identify the incorrectly predicted sessions where the ranking list fails to hit the ground truth item at time step $l+1$ (i.e., $v_{l+1}^{s}\notin O(s)\O$). Afterward, we prompt the LLM to conduct self-reflection, that is, analyze and summarize the prediction errors, to generate potential hints in natural language to rectify such errors. These hints thus can be considered as specialized knowledge for better recommendations. Inspired by Chain-of-Thought prompting (Wei et al., 2022), our approach focuses on analyzing and alleviating these errors in a step-by-step and progressive reasoning chain, producing hints that are tailored to be compatible with LLM comprehension capabilities, as introduced below.
@@ 3.1.2. Automated Filtering
To ensure the effectiveness, we only add a new hint to $\mathcal{H}$ if it can lead to performance improvement. Specifically, we first construct a hint-enhanced prompt $P^{*}$ (Prompt 5) to trigger the LLM with the hint-enhanced prompt for recommendation inference:
(2) $O^{*}(s)=LLM(P^{*}(s,\mathcal{C},h)),$
where $O^{*}(s)$ is the recommendation list for session $s$ based on the hint-enhanced prompt $P^{*}$. Then, we compare $O^{*}(s)$ with $O(s)$ obtained by Eq. 1 using the basic prompt without hint enhancement. The hint $h$ can be recognized as effective if it leads to a better prediction by the LLM, i.e., $v_{l+1}^{s}\in O^{*}(s)\And v_{l+1}^{s}\notin O(s)$.
Prompt 6
Meanwhile, some hints may be redundant generated by LLMs, as multiple sessions may suffer from similar causes. For example, ‘Consider the release years of movies in the user’s history for era preference’ and ‘Think about the production year of watched movies, recommend movies in that era from the candidate set’ share a similar semantic meaning. The redundancy of hints may result in a high cost of the maintenance and utilization of the hint knowledge base $\mathcal{H}$, e.g., larger size of the hint knowledge base and increased complexity in retrieving the targeted hint. To reduce the redundancy in the hint knowledge base $\mathcal{H}$, we only incorporate new hints that are distinct from existing ones. Specifically, we employ LLMs to detect the semantic similarity between the candidate hint $h^{\prime}$ and existing ones $\{h\in\mathcal{H}\}$, given by,
@@ 3.2. Reinforcement Utilization Module
To guide LLMs to infer more accurate SBR predictions, we propose to utilize the constructed hint knowledge base to prevent errors in the inference process of LLMs.
In practical scenarios, the absence of explicit labels on the hints’ efficacy leads to a challenge in conducting supervision on hint selection, as computing the efficacy for all hints on each sample is costly and redundant. Fortunately, DRL with a replay buffer enables us to collect reward signals (i.e., hint efficacy) and update the network spontaneously to speed up. To this end, we employ the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm to ensure stable and efficient training of our retrieval agent. Specifically, our agent is trained to select the most relevant hints based on session-related context information, thereby preventing LLMs from similar reasoning mistakes for recommendation. In addition, the proposed DRL framework allows for the balance between exploitation and exploration. It is also compatible with more complex extensions such as retrieving multiple hints and multi-round agent-based interactions.
@@ 3.2.1. Markov Decision Process (MDP)
Transition. In each step, the agent can receive task-specific feedback by simulating the real-world RSs. The agent observes the state of the current session, takes action to select a hint for the LLM, and receives a reward from the environment. Subsequently, the environment transits to the next session’s state, and then the transition is denoted as $T(\mathbf{z})=\mathbf{z}^{\prime}$. This setup can flexibly accommodate the scenario where multiple steps per session are involved for more sophisticated learning processes in our future studies.
Replay buffer. To facilitate efficiency in policy optimization, we maintain a replay buffer $D=(\mathbf{z},\mathbf{a},r,\mathbf{z}^{\prime})$ to store the tuples of observation state, agent action, reward, and next observation state. With the records in the replay buffer, the retrieval agent can refine successful strategies and learn from erroneous trials. This technique significantly accelerates the DRL training as the LLM backbone is relatively slow for inference and producing reward signals.
@@ 3.2.2. PPO Training
where the first term is the reward measured by recommendation tasks, and the second term is the KL-divergence of policies with a coefficient $\beta$ for the regulation in policy updates. In summary, we train a retrieval agent with task-specific feedback via DRL, which can select the appropriate action to improve LLM’s performance through hint-enhanced prompts for better recommendations.
Input: Training data $\mathcal{S}_{tr}$ for SBR, hint knowledge base size $n_{h}$.
@@
Input: Training data $\mathcal{S}_{tr}$ for SBR, constructed $\mathcal{H}$, max episode $N$.
@@
3.2.3. Retrieval-Enhanced LLMs for Recommendation
With the constructed knowledge base and the trained retrieval agent, we achieve more accurate recommendations by automatically selecting appropriate hints for prompt enhancement during inference. The final recommendation output $\hat{O}$ can be denoted as follows:
(6) $\hat{O}=LLM(P^{*}(s,\mathcal{C},\pi_{\bm{\theta}}(\mathbf{z}_{s}))),$
@@ 4.1.2. Baselines
We compare our proposed method with eight state-of-the-art baseline methods [^5] Our method does not require fine-tuning, and due to the high computational cost, fine-tuning-based methods are not in our scope. , and the details are listed as follows. FPMC (Rendle et al., 2010) combines matrix factorization with the first-order Markov chain. GRU4Rec (Hidasi et al., [n. d.]) uses Gated Recurrent Unit (GRU), a typical RNN layer, to model whole sessions. NARM (Li et al., 2017) introduces the attention mechanism into RNNs to better model the sequential behavior. SRGNN (Wu et al., 2019) constructs session graphs and employs GNN layers to obtain embeddings. GCEGNN (Wang et al., 2020) further includes global-level information as an additional graph to enhance the model performance. AttenMixer (Zhang et al., 2023b) achieves multi-level reasoning over item transitions by the attention-based readout method. LLMRank (Hou et al., 2023) is the LLM-based method, which demonstrates the promising zero-shot inference ability of LLMs through recency-focused prompting and in-context learning. NIR (Wang and Lim, 2023) is also LLM-based method that employs a multi-step prompt to capture user preferences first, then select representative movies, and finally perform the next-item prediction.
@@
4.1.3. Evaluation Strategy and Metrics
For each dataset, we apply the split-by-ratio strategy (Sun et al., 2020) for sessions to obtain training, validation, and test sets with a ratio of 7:1:2. For the full dataset setting, we use the entire training set. For the few-shot setting, we sample 500 training samples from the entire training set. In terms of evaluation, we consider the last item in each session as the target. Considering the cost and efficiency of LLMs, we randomly sample a subset of 2,000 sessions from the test set. For the efficiency of evaluation, we sampled 49 negative items for each positive item (i.e., target). In addition, we adopt two commonly used metrics for model evaluation, including normalized discounted cumulative gain ($NDCG@n$) and hit rate ($HR@n$), with $n=\{5,10\}$. The results show the average of five runs with different random seeds.
4.1.4. Implementation Details
For a fair comparison, all methods are optimized by the Adam optimizer with the same embedding dimension $768$ that aligns with the BERT encoder dimension [^7] Empirically, using MLP to reduce dimension leads to a consistent performance drop.. Following (Sun et al., 2022), we employ the Optuna [^8] https://optuna.org/ framework to efficiently optimize the hyper-parameter of all methods. We conduct 20 trials on the following searching spaces, i.e., learning rate: {$1e^{-2},1e^{-3},1e^{-4}$}, weight decay: {$1e^{-4},1e^{-6}$ $,1e^{-8}$}, and batch size: {$16,64,256$}. For our method Re2LLM, we set candidate set size $|\mathcal{C}|$ to 50, knowledge base size $n_{h}$ to 20, and few-shot training size $n_{e}$ to 500. The impacts of essential hyper-parameters on Re2LLM can be found in Section 4.4. For LLM-based methods, we use ‘gpt-4’ API as the backbone model for these methods. For the rest hyper-parameters in baseline methods, we follow the suggested settings in the original papers. The experiments are run on a single Tesla V100-PCIE-32GB GPU. On average, around 50 data samples are required to generate 20 hints. The average cost to analyze and conduct inference on each sample is around USD 0.015 during the hint generation, DRL training, and final testing.
@@ 4.3. Ablation Studies (RQ2)
Table 3 shows the performance of all ablation studies. First, Re2LLM outperforms its variant w/o-HE, which indicates the necessity of activating LLM’s reflection mechanism with hint enhancement in our Reflective Exploration Module. Second, Re2LLM outperforms variants w/o-DRL(RAN) and w/o-DRL(ALL), indicating the importance of our Reinforcement Utilization Module for the training of the retrieval agent with task-specific feedback signals. The variant w/o-DRL(RAN) has the worst performance by random sampling hints because inaccurate hints can mislead the LLM inference for recommendation, resulting in negative impacts. The variant w/o-DRL(ALL) shows limited improvement compared to the variant w/o-HE, revealing that simply feeding all hints into LLMs without session-aware selection is not optimal. Third, the variant w/o-AF achieves relatively better performance among all variants. Nevertheless, it still underperforms our Re2LLM due to the positive impact of the automated filtering mechanism on the hint knowledge base construction. In summary, the above results exhibit the efficacy of different modules in the proposed Re2LLM for more accurate SBR.
4.4. Hyper-parameter Analysis (RQ3)
4.4.1. Knowledge Base Size
We investigate the correlation between model performance and the size of the knowledge base constructed by the Reflective Exploration Module. As shown in Fig. 3, there is an improvement in both evaluation metrics as the size of the knowledge base increases to 20. The improvement is attributed to the diversity of the generated reflective hints by covering a broader range of common errors of LLM inference on SBR. However, when the knowledge base size becomes too large, there is a slight performance drop due to the increased complexity and difficulty in the retrieval agent optimization. Thus, we suggest adopting a moderate hint knowledge base size.
Figure 4. Performance of the representative baselines and our method with varying few-shot training size for retrieval agent. The dotted lines indicate w/o-DRL(RAN) performance.
4.4.2. Few-shot Training Size.
We also study how model performance is affected by the number of few-shot training samples used for the retrieval model training. Fig. 4 shows the comparison of Re2LLM, its variant w/o-DRL(RAN), and representative baselines across various few-shot training sizes [^9] We use previously mentioned data augmentation (Tan et al., 2016) for baselines. The results are from the Movie domain. Similar trends occur for the VideoGame domain.. First, using too few samples (e.g., less than 50) for retrieval agent training results in inferior performance, which is even lower than w/o-DRL(RAN) with random hint selection. This is because the retrieval agent is under-trained with only very few samples. Second, as the number of training samples increases, the performance of our method increases consistently, showing the growth of the generalization ability of the retrieval agent. Baseline methods also improve, but remain inferior to our proposed method across various few-shot training sizes. Third, we also observe the limited improvement of Re2LLM when the number of samples exceeds a threshold. For the balance between effectiveness and efficiency, 500 is an appropriate selection for the few-shot training size of Re2LLM.
4.4.3. Reward Design
Table 4. Performance of Re2LLM with two reward designs.
Reward Movie Game
HR@10 NDCG@10 HR@10 NDCG@10
absolute 0.5143 0.3147 0.7026 0.4542
comparative 0.5735 0.3358 0.7173 0.4689